OpenStudy (anonymous):

Savannah is 24 years older than Ryan, while Sebastian is 5 times as old as Ryan. If Savannah and Sebastian are twins, which equation may be used to determine the age of all three people?

jhonyy9 (jhonyy9):

let Savanna note x Ryan note y Sebastian note z x=24+y z=5y x=z this substitute in place of x in the first equation and will get one system of equations z=24+y z=5y this just you need substitute in place of z in the first equation and will get the value of y - using this value of y you can calculi easy the value of z and in this way because z=x you have solved this system for x,y and z hope these will help you sure